OpenStudy (anonymous):

5sqrt{3}-sqrt{98}+sqrt{48}+6sqrt{2}

OpenStudy (mathstudent55):

Is this it? \(5\sqrt{3}-\sqrt{98}+\sqrt{48}+6\sqrt{2}\)

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

The first and last roots are roots of prime numbers. They can't be simplified. You need to simplify the roots of 98 and 48.

OpenStudy (mathstudent55):

Look at 98 first. Can you list the prime factors of 98?

OpenStudy (anonymous):

is it simplified to 9sqrt3-sqrt2

OpenStudy (mathstudent55):

\(5\sqrt{3}-\sqrt{98}+\sqrt{48}+6\sqrt{2}\) \(= 5\sqrt{3}-7\sqrt{2}+4\sqrt{3}+6\sqrt{2} \) \(= 9\sqrt{3}-\sqrt{2} \) You are correct.